(a) Find the linearization of f(x, y) = e^(x) cos y at the point (0, π/2)



1
Expert's answer
2021-08-09T10:12:44-0400

Solution;

The linearization of a function f(x,y) at a point (x0,y0) is given by;

"L(x,y)=f(x_0,y_0)+( x-x_0)f_x(x_0,y_0)+(y-y_0)f_y(x_0,y_0)"

Given;

"f(x,y)=e^xcos y" and "(x_0,y_0)=(0,\fracπ2)"

"f(0,\fracπ2)=e^0cos(\fracπ2)=0"

We can the derive;

"f_x=e^xcos y"

"f_y=-e^xsiny"

From which ;

"f_x(0,\fracπ2)=e^0cos(\fracπ2)=0"

"f_y(0,\fracπ2)=-e^0sin(\fracπ2)=-1"

Hence by substitution;

"L(x,y)=0+0(x-0)+(y-\fracπ2)(-1)=\fracπ2-y"

"L(x,y)=\fracπ2-y"